Four point masses of mass each of 1 kg are connected by a string which forms a square of diagonal 0.707 m . If square is placed on rotating table rotating with 5 rotation per second then what is the tension in string ?


1
Expert's answer
2021-01-25T13:56:04-0500

Answer

Given diagonal =0.707 meter

Lets take length of square is a meter diagonal will be

"\sqrt{2}" a meter


"\sqrt{2}a" =0.707 meter

a=0.5 meter


Rotation is w=5 rps

=5×2π=10πrad/s

We all know that

Radius for revolution = half of diagonal

"=\frac{a}{\sqrt{2}}"

So we can write

"2T\cos 45^°=mw^2r"

Putting value of m, w, r then the tension

T=246.8N